# Diffusion approximation of the stationary distribution of a two-level single server queue

### v6

@@ 5.1 Process limit
Thus, (A.34) and (A.35) prove (4.24) for $k=1,2$. Similarly, it follows from (4.22) with $B=(1/r,\infty)$ that, for $k=1,2$,
$|\mathbb{E}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}[(R^{(r)}_{e})^{k}]|$ $\leq\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}>1/r)]$
$=\frac{1}{k+1}\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}1(T^{(r)}_{e,i}>1/r)]=o(r^{2-k}),$
To prove (4.26), we apply $f({\boldsymbol{x}})=x_{2}^{k}1(x_{2}>1/r,x_{3}>0)$ to (4.1), then
$k\mathbb{E}[(R^{(r)}_{e})^{k-1}1(R^{(r)}_{e}>1/r)]$ $=\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}1(R^{(r)}_{e}>1/r)]$
$\quad+\sum_{i=1,2}\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,i})^{k}1(T^{(r)}_{e,i}>1/r)]\mathbb{P}^{(r)}_{e}[w_{i}(L^{(r)}(0-))]$
$\geq\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}1(R^{(r)}_{e}>1/r)].$
Hence, applying (4.22) with $B=(1/r,\infty)$ and $k-1$ instead of $k$ to the above formula, we have
$|\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e})^{k}]|$ $\leq\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}1(R^{(r)}_{e}>1/r)]$
$\leq(\alpha^{(r)}_{e})^{-1}k\mathbb{E}[(R^{(r)}_{e}(0))^{k-1}1(R^{(r)}_{e}>1/r)]$
$\leq k\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k}1(T^{(r)}_{e,i}>1/r)],$
where the last term is $o(r^{2})$ for $k=1$ and $o(r)$ for $k=2$ by (A.34) and (A.35). Thus, we have (4.26). (4.27) is similarly obtained.
@@
From Taylor expansion of twice differentiable function $f$:
$|f(x)-(f(0)+f^{\prime}(0)x)|\leq\frac{1}{2}|x|^{2}\max_{|u|\leq|x|}|f^{\prime\prime}(u)|,\qquad x\in\mathbb{R},$ (A.36)
we have $|e^{-x}-1+x|\leq\frac{1}{2}|x|^{2}e^{|x|}$. Hence, for $i=1,2$,
$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}(\ell_{0}+o(r))}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}o(r)),$
$e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}(\ell_{0}+r\theta_{i}+o(r))}=e^{r\theta_{i}\ell^{(r)}_{0}}(1+\theta_{i}r+\theta_{i}^{2}o(r)).$
Thus, (4.35) is obtained. Using these asymptotic expansions, we have
$e^{r\theta_{2}(\ell^{(r)}_{0}+1)}-e^{r\theta_{1}(\ell^{(r)}_{0}+1)}$
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}}+e^{r\theta^{(r)}_{2}\ell^{(r)}_{0}}(\theta_{2}r+\theta_{2}^{2}o(r))-e^{r\theta^{(r)}_{1}\ell^{(r)}_{0}}(\theta_{1}r+\theta_{1}^{2}o(r))$
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}}+r(e^{r\theta^{(r)}_{2}\ell^{(r)}_{0}}\theta_{2}-e^{r\theta^{(r)}_{1}\ell^{(r)}_{0}}\theta_{1})+(e^{r\theta_{2}\ell^{(r)}_{0}}\theta_{2}^{2}-e^{r\theta_{1}\ell^{(r)}_{0}}\theta_{1}^{2})o(r)$
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}}+r(e^{\theta_{2}\ell_{0}}(1+\theta_{2}o(r))\theta_{2}-e^{\theta_{1}\ell_{0}}(1+\theta_{1}o(r))\theta_{1})$
$\qquad+(e^{r\theta_{2}\ell^{(r)}_{0}}\theta_{2}^{2}-e^{r\theta_{1}\ell^{(r)}_{0}}\theta_{1}^{2})o(r)$
$\quad=e^{r\theta_{2}\ell_{0}}-e^{r\theta_{1}\ell_{0}}-r(e^{\theta_{2}\ell_{0}}\theta_{2}-e^{\theta_{1}\ell_{0}}\theta_{1})+(e^{r\theta_{2}\ell^{(r)}_{0}}\theta_{2}^{2}+e^{r\theta_{1}\ell^{(r)}_{0}}\theta_{1}^{2})o(r).$
This proves (4.36), and (4.37) is similarly proved. ∎
@@
It follows from Lemma 4.2 that
$e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}=1-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}+\frac{1}{2}|\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}|^{2}e^{|\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}|}$
$\quad=1+(\mu r\theta_{2}+O(r^{2}))\widehat{R}^{(r)}_{s}+O(r^{2})(\widehat{R}^{(r)}_{s})^{2}$
$\quad=1+\mu r\theta_{2}\widehat{R}^{(r)}_{s}+O(r^{2})(\widehat{R}^{(r)}_{s})^{2}.$
This proves (4.39), and (4.8) and (4.8) are similarly proved. ∎
@@ Acknowledgement
I thank the editor and referees for their helpful comments and careful reading. This study is motivated by the BAR approach, coined by Jim Dai (e.g., see [3]). I am grateful to him for his grand design about the BAR approach, and to Rami Atar for providing his recent paper [1]. I also benefit from Evsey Morozov for his comments on this paper.

### v7

@@ 6.1 Process limit
Thus, (4.28) is proved for $k=1,2$. We next prove (4.29) for $v=e$. Since $\mathbb{E}[(R^{(r)}_{e})^{k}]$ is uniformly bounded for $k=1,2$ by (4.27), we can prove (4.29) by the same argument as the proof of (4.28), but the order of $r$ is one less because of (4.27). We omit the proof for $v=s$ because it is similar to the case of $v=e$ except for (4.26) to be used instead of (4.27).
@@
For a $(n+1)$-times continuously differentiable function $f$, we have, from Taylor expansion (e.g. see Theorem 19.9 of [14]),
[…]
$\leq\frac{1}{(n+1)!}|x|^{n+1}\max_{|u|\leq|x|}|f^{(n+1)}(u)|,\quad x\in\mathbb{R},n\geq 0.$ (A.36)
Applying $f(x)=e^{-x}$ to this inequality, we have $|e^{-x}-1+x|\leq\frac{1}{2}|x|^{2}e^{|x|}$. Hence, for each fixed $\theta\in\mathbb{R}$, as $r\downarrow 0$
$e^{r\theta\ell^{(r)}_{0}}=e^{\theta(\ell_{0}+o(r))}=e^{\theta\ell_{0}}(1+\theta o(r))+o(\theta o(r))=e^{\theta\ell_{0}}(1+o(\theta r)),$
$e^{r\theta(\ell^{(r)}_{0}+1)}=e^{\theta r\ell^{(r)}_{0}}e^{r\theta}=e^{\theta\ell_{0}}(1+o(\theta r))(1+\theta r+o((\theta r)^{2}))=e^{\theta\ell_{0}}(1+\theta r+o(\theta r)).$
Thus, (4.37) is obtained. (4.38) and (4.7) are immediate from these asymptotic expansions. ∎
@@
Since $\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}$ and $\eta^{(r)}_{i}(r\theta)\widehat{R}^{(r)}_{e}$ are uniformly bounded in $r<a$ for a sufficiently small $a>0$ and vanish as $r\downarrow 0$, it follows from Lemma 4.2 and (A.4) that
$e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}=1-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}+\frac{1}{2}(\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-})^{2}+O\big{(}|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|^{3}\max_{r|\theta|<a}e^{|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|}\big{)}$
$\quad=1-\left(-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}(r\theta)^{2}+o((r\theta)^{2})\right)\widehat{R}^{(r)}_{s-}$
$\qquad+\frac{1}{2}\left(-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}(r\theta)^{2}+o((r\theta)^{2})\right)^{2}(\widehat{R}^{(r)}_{s-})^{2}$
$\qquad+(\widehat{R}^{(r)}_{s-})^{3}O\big{(}(r\theta)^{3}\max_{r|\theta|<a}e^{|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|}\big{)}$
$\quad=1+r\theta(\mu^{(r)}\widehat{R}^{(r)}_{s-})-\frac{1}{2}(r\theta)^{2}\left((\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}\right)+S_{3}(\widehat{R}^{(r)}_{s-})o((r\theta)^{2}),$
where recall that $S_{3}(\widehat{R}^{(r)}_{s-})=\widehat{R}^{(r)}_{s-}+(\widehat{R}^{(r)}_{s-})^{2}+(\widehat{R}^{(r)}_{s-})^{3}$. This proves (4.8), and (4.8) and (4.8) are similarly proved. ∎
A.5 Proofs of Lemmas 5.1 and 5.2
We first prove Lemma 5.1, namely, that (5.3) implies (5.4). To prove this, we use the BAR (4.1) for $f(z,x_{1},x_{2})=(x_{1}\wedge 1/r)(x_{2}\wedge 1/r)1(z=\ell^{(r)}_{0})$, then we have
$\mathbb{E}[R^{(r)}_{e}1(R^{(r)}_{e}\leq 1/r,L^{(r)}=\ell^{(r)}_{0})]+\mathbb{E}[R^{(r)}_{s}1(R^{(r)}_{s}\leq 1/r,L^{(r)}=\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}\mathbb{E}[\widehat{T}^{(r)}_{e,2}]\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\alpha^{(r)}_{e}\mathbb{E}[\widehat{T}^{(r)}_{s}]\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (A.37)
Then, we can see that the left-hand side of this equation vanishes as $r\downarrow 0$ by (5.3) because, by Schwartz’s inequality,
$\big{(}\mathbb{E}[R^{(r)}_{v}1(R^{(r)}_{v}\leq 1/r,L^{(r)}=\ell^{(r)}_{0})]\big{)}^{2}\leq\mathbb{E}[(R^{(r)}_{v})^{2}]\mathbb{P}[L^{(r)}=\ell^{(r)}_{0}],\qquad v=e,s.$
Hence, (5.4) follows from (A.5) because $\mathbb{E}[\widehat{T}^{(r)}_{s}]$ converges to $\mu^{-1}$ as $r\downarrow 0$. Thus, Lemma 5.1 is proved.
We next prove Lemma 5.2. By Schwartz’s inequality,
$\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0}]\big{\}}^{2}=\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3/2}(\widehat{R}^{(r)}_{e})^{1/2}1^{2}(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}^{2}$
$\quad\leq\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{0})]\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})].$
Since $\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ is uniformly bounded in $r$ by Corollary 4.4, (5.4) implies (5.5). This complete the proof of Lemma 5.2.
@@ Acknowledgement
This study is motivated by the BAR approach, coined by Jim Dai (e.g., see [4]). I am grateful to him for his grand design about it, and to Rami Atar for providing his recent papers [1, 2]. I also benefit from Evsey Morozov for his comments on this paper.